1/1

Manfred Hausmann, Manfred Hausmann,

Manfred Hausmann,

Fēngqiáo Yè Pō Maple Bridge Night Mooring Maple Bridge Night Mooring

tradukita de Jì Zhāng tradukita de Zhao Yanchun

tradukita de Hugh Grigg

Yuè luò wū tí shuāng mǎn tiān,

jiāng fēng yú huǒ

duì chóumián. Gūsū chéng wài

hán shānsì, yèbàn zhōng shēng

dào kèchuán.

The moon setting, crows cawing, frost overfilling the skies, The maples riv’rside, lamps aboard, sad the fishermen’s eyes.

The toll from the Cold Mountain Temple outside of Suchow Reaches my boat at midnight and floods my sleepless sorrow.

The moon descends, crows caw, frosti-

ness fills the sky; riverside maples, fishermen’s lights, facing me in my worried sleep. Ouside Gusu city lies Cold Mountain

Temple; at midnight, the sound of bells reaches

the ferries.

Traduko de la Germana poemo “An der Ahornbrücke” de Manfred Hausmann (∗1898-09-10 – †1986-08-06) en ĉina de Jì Zhāng .

Arg-911-1819 (2013-08-19 12:33:18)

Vidu ankaŭ: http: // eastasiastudent. net/ china/ wenyan/ zhang-ji-maple-bridge/ .

Traduko de la Germana poemo “An der Ahornbrücke” de Manfred Hausmann (∗1898-09-10 – †1986-08-06) en la Anglan de Zhao Yanchun.

Arg-911-1824 (2013-07-27 10:23:29)

Traduko de la Germana poemo “An der Ahornbrücke” de Manfred Hausmann (∗1898-09-10 – †1986-08-06) en la Anglan de Hugh Grigg.

Arg-911-1823 (2013-07-27 09:27:49)

Pri la tradukinto vidu la retejon http: // eastasiastudent. net/ about/ .

http://www.poezio.net/version?poem-id=911&version-id=1819,1824,1823